# Observation of diamagnetic strange-metal phase in sulfur-copper codoped lead apatite

### v1

@@ Abstract
By codoping sulfur and copper into lead apatite, the crystal grains are directionally stacked and the room-temperature resistivity is reduced from insulating to $2\times 10^{-5}~{}\Omega\cdot$m. The resistance-temperature curve exhibits a nearly linear relationship suggesting the presence of strange-metal phase. A possible Meissner effect is observed in dc magnetic measurements. An even-in-field transverse voltage is found indicating the possible contribution of individual vortex dynamics. A clear pathway towards superconductivity in this material is subsequently benchmarked.
@@
Materials with multiple elements, such as cuprate superconductor, make diversity in physics. It is always difficult to deal with multiple elements, since each of them is individually activated at distinct temperature. Copper-substituted lead apatite, also named as LK-99 Lee et al. (2023a, b), was claimed to be room-temperature superconductor, but due to the complicated components and structures, the reproduction is still controversial Wang et al. (2023); Guo et al. (2023); Habamahoro et al. (2023); Liu et al. (2023a, b); Wang et al. (2024). Considering copper is a sulfophilic element, the participation of sulfur in the synthesis benefits the substitution of copper, which does originally not favor the ionic crystal structure of lead apatite. As reported in our previous work Wang et al. (2024), this sulfur-copper codoped lead apatite (SCCLA) manifests a weak Meissner effect at near room temperature. In order to further enhance the effect, we have to finely optimize the reaction procedure as sulfur could not be held in the bulk at overhigh temperature that enables other elements to react. In this work, we then adjust the synthetic procedure of SCCLA and find the signal magnitude is largely increased.
@@
Figure 1: (a) Schematic of sample synthesis procedure of sulfur doping. (b) XRD pattern of the sample after 24 hours soaking in sulfide solution. (c) The morphological characteristics with the hexagonal grains highlighted. (d) Schematic of the directional stacking mechanism under high pressure.
@@
Pure copper-substituted lead apatite was first synthesized through a hydrothermal-calcination method following the procedure outlined in the literature Lee et al. (2023a); Wang et al. (2024), with the resulting gray powder being insulating. In order to carry out sulfur doping, as sketched in Fig. 1(a), the powder is ball-milled to nano-scaled particles and then thoroughly dispersed in a saturated solution of sodium sulfide. After stirring for 12 hours, the sample undergoes a 24-hour reaction in a high-pressure reaction vessel. The internal pressure of the reaction vessel is maintained at no less than 4 MPa, and the reaction temperature is set to 170 ${}^{\circ}$. After the reaction, the powder is washed, filtered, dried, and then subject to a pressure of 20 MPa to compress it into a solid tablet.
@@
The XRD spectra of the sample are refined using Reflex as shown in Fig. 1(b). Pawley Fitting is utilized to accurately determine lattice constants, and experimental and sample-related parameters. During the sulfur doping process, atoms within the apatite structure may have been etched or consumed, so the lattice is largely shrunk leading to a strong lattice distortion and good conductivity. The experimental spectra are consistent with that of the phosphate of P63m structure. By comparing to the standard of lead phosphate, the intensity differences observed at some peak positions could be attributed to the orientation of nano-scaled grains during growth in the sulfide solution. SEM was utilized to observe the morphological characteristics of the sample, as displayed in Fig. 1(c). The powder mainly consists of hexagonal flakes or hexagonal crystals. A small amount of hexagonal prism rod-like particle shows up, pointing out a significant difference compared to the sample before sulfur doping. This observation together with the analysis results of XRD spectra reveal the microstructure of the powder emerging as hexagonal grains, so the directional stacking is evident to function during high-pressure agglomeration, as sketched in Fig. 1(d). The head-to-tail quasi-1D stacking structure is then changed to be side-by-side, so the conducting channel might be broadened.
@@
Figure 2: (a) MT curves of ZFC and FC measurements at 25 and 200 Oe, respectively. The signal of an empty capsule is shown for comparisons. (b)-(e) MH hysteresis curves at eight temperatures. At 10 and 50 K, it is paramagnetic at strong field, and at higher temperature it becomes diamagnetic. Below 250 K, a remarkable magnetic hysteresis effect is observed between $-300$ and 300 Oe. At 300 and 350 K, although we have decreased the measurement interval, there is not visible hysteresis indicating the critical temperature is below 300 K. (f) MH hysteresis curve at 250 K is again finely measured with a quartz holder.
@@
We then employed MPMS-3 SQUID to measure the dc magnetic moments of the synthesized SCCLA samples, as shown in Fig. 2. It is clear from the magnetic moment–temperature (MT) curves, considering the background of capsule, the magnetization under magnetic field of 25 and 200 Oe is negative up to room temperature, indicating its diamagnetism at weak field. The bifurcation between zero-field-cooling (ZFC) and field-cooling (FC) measurements is also found to be around room temperature. The moment–magnetic field (MH) curves at various temperatures from 10 to 350 K are also displayed. One can see that, from $-300$ to 300 Oe, the MH curves exhibit notable hysteresis effect up to 250 K. In particular, the hysteresis at 150 K is pretty obvious, exceeding the highest critical temperature of known superconductors at ambient pressure. This phenomenon has been reported in our previous paper Wang et al. (2024), but the quality of the present data is largely increased, further eliminating the possibility of measurement faults. At 300 and 350 K, the hysteresis can not be detected. As stated before, these results strongly suggest there exists a Meissner effect in SCCLA.
@@
Figure 3: (a) The RT curves and their derivative at current of 8 mA during heating and cooling the sample. Based on the RT relationship, the temperature region can be divided into three: Fermi liquid, strange metal and bad metal. (b) The IV curves at four temperatures are displayed.
@@
Measurement results of transport are displayed in Fig. 3, which are mainly conducted on PPMS and have been reconfirmed on other facilities. Although the sample before soaking in sulfur is insulator, after soaking the resistivity at room temperature is largely reduced to around $2\times 10^{-5}~{}\Omega\cdot$m, close to that of natural graphite. It is thus evident that the sulfur-copper codoping plays essential role in the improvement of transport properties of the insulating apatite ionic crystal. It is also worth noting that, the tablet samples are just mechanically compressed from nano-scaled powders which is too fragile to fix electrodes, so we believe there is still large room for further improving the conductivity. Nonetheless, such considerable conductivity is still far beyond expectation, that has to be carefully comprehended.
@@
As estimated by the critical magnetic field, if there is a superconducting phase, the critical current should be as small as $\mu$A, which can not be accurately detected in our present facility. The RT curves are therefore measured with constant current of 8 mA to ensure the precision, as illustrated in Fig. 3(a). It is found that below 20 K the derivative $dR/dT$ exhibits a nearly linear relation suggesting a Fermi liquid behavior as in normal metals. Increasing the temperature, the derivative becomes nearly constant revealing a linear-in-temperature resistivity, which is an essential signature of strange metal. At around 230 K in the cooling curve and 304K in the heating curve, there are remarkable turning points to figure out the transition from strange metal to bad metal phase. It has been stated that, if the resistivity in bad metal can be further reduced by one or two orders Yang et al. (2022), these turning points can then be recognized as the superconducting critical temperature. The thermal hysteresis of the two RT curves, probably stemming from the poor thermal conductivity of the sample, is also consistent with that of the magnetic measurements. The current–voltage (IV) curves at various temperatures are displayed in Fig. 3(b). The step of measurement is chosen to be 400 $\mu$A, so the curves are all linear obeying Ohm’s law. As an additional note, during the first test of signal-noise-ratio in IV curve which has to be reconfirmed in the future, we set a relatively small step and find the voltage does not increase within 10 mA. A possible superconducting phase might be fragile when large current is injected.
Figure 4: Hall resistance (a) and longitudinal magnetoresistance (b) at four temperatures and current of 8 mA. The even part (c) and odd part (d) of Hall resistance are obtained by $(R_{xy,+}+R_{xy,-})/2$ and $(R_{xy,+}-R_{xy,-})/2$, respectively, with $R_{xy,\pm}$ being the Hall resistance at positive or negative magnetic field.
One now would be wondering why the strange metal exhibits a possible Meissner effect. We consider the diamagnetic grains are nano-scaled and stacked so the vortex should cross the margins to result in resistance of strange metal. To this end, we detect the Hall ($R_{xy}$) and longitudinal ($R_{xx}$) resistance, i.e. magnetoresistance (MR), at various temperatures as shown in Fig. 4. At 10 and 300 K, both Hall and MR are around $0.5\%$, slightly larger than those at other temperatures. It is noticed that the Hall resistance is not centrosymmetric as in the normal case. We thus obtain its even and odd part by calculating $(R_{xy,+}+R_{xy,-})/2$ and $(R_{xy,+}-R_{xy,-})/2$, with $R_{xy,\pm}$ being the Hall resistance at positive or negative magnetic field. The even part of Hall resistance draws our attention which is larger than the odd part. Normally speaking, the even part often stems from the longitudinal MR as the Hall bar can not be ideally aligned. Herein, however, we notice that the MR itself does not manifest an evenly symmetric lineshape at high temperatures. Hence, the appearance of the even-in-field transverse voltage (ETV) seemingly reflects the intrinsic features of the sample da Luz et al. (2005); Guryev et al. (2019); Baumans et al. (2016); Francavilla et al. (1995); Segal et al. (2011); Francavilla and Hein (1991). Different from the Lorentz force induced normal Hall effect, which gets inverse sign by reversing magnetic field, the ETV is normally regarded to be from the vortex creep dynamics across margins of phases. The magnetic field mainly influences the hopping rate of creep but not the direction. In our opinion, this perspective of individual vortex dynamics can thus be the most reasonable interpretation for both the magnetic and electric properties.
@@
Compared to our previous synthetic procedure Wang et al. (2024), the doping of sulfur not only facilitates the doping of copper, but also changes the topology of 1D ionic channel of apatite. From head-to-tail to side-by-side, the quasi-1D copper-sulfur lattice may have got much stronger inter-chain interaction to activate tunneling in between. This explains the huge improvement of electric conductivity and the emergence of strange metal. We have also conducted microwave absorption measurement and did not find visible radical signals, implying the absence of quasiparticles in normal metal. People might suspect the magnetic hysteresis is enabled by some pseudogap states, as we do not explicitly observe a zero-resistance state along with the possible Meissner effect. But since the RT curves definitely exhibit a metallic feature, we would be more preferable to think it is because the superconducting critical current is below our measurement limit. During measurements, we have noticed that the sample possesses abnormally large electric capacity and poor thermal conductivity Bi et al. (2022); Deppe et al. (2004), which are inconsistent with the high electric conductivity. In particular, the capacity often induces strange sudden change between low- and high-resistance states, even if we have carefully excluded other interferences. It is more likely the sample is continuously charging and discharging during electric measurements, hindering the detectability of zero resistance. The dielectrics of lead apatite backbone matters.
@@
In summary, we modify the synthetic procedure of SCCLA to codope both sulfur and copper into lead apatite, and the structural characterization reveals a directional stacking mechanism. The magnetic and electric properties of SCCLA have been comprehensively investigated. The hysteresis MH loops can be observed up to 250 K, and the ZFC–FC bifurcation is around 300 K. The RT curve manifests that SCCLA possesses a strange-metal phase at large current. A weak even-in-field transverse voltage indicates the possible contribution of superconducing vortex dynamics. We therefore believe that we have made a substantial step towards room-temperature superconductivity.

### v3

@@ Abstract
By codoping sulfur and copper into lead apatite, the crystal grains are directionally stacked and the room-temperature resistivity is reduced from insulating to $2\times 10^{-5}~{}\Omega\cdot$m. The resistance-temperature curve exhibits a nearly linear relationship at low temperature suggesting the presence of strange-metal phase, and a second-order phase transition is then observed at around 230 K during cooling the samples. A possible Meissner effect is present in dc magnetic measurements. Further hydrothermal lead-free synthesis results in smaller resistance and stronger diamagnetism, demonstrating the essential component might be sulfur-substituted copper apatite and the alkalis matter as well. A clear pathway towards superconductivity in this material is subsequently benchmarked.
@@
Materials with multiple elements, such as cuprate superconductor, make diversity in physics. It is always difficult to deal with multiple elements, since each of them is individually activated at distinct temperature. Copper-substituted lead apatite, also named as LK-99 Lee et al. (2023a, b), was claimed to be room-temperature superconductor, but due to the complicated components and structures, the reproduction is still controversial Wang et al. (2023); Guo et al. (2023); Habamahoro et al. (2023); Liu et al. (2023a, b); Wang et al. (2024). Considering copper is a sulfophilic element, the participation of sulfur in the synthesis benefits the substitution of copper, which does originally not favor the ionic crystal structure of lead apatite. As reported in our previous work Wang et al. (2024), this sulfur-copper codoped lead apatite (SCCLA) manifests a weak Meissner effect at near room temperature. In order to further enhance the effect, we have to either finely optimize the reaction procedure as sulfur could not be held in the bulk at overhigh temperature that enables other elements to react, or alternatively design and synthesize some lead-free frameworks. In this work, we then adjust the synthetic procedure of SCCLA and find the signal magnitude is largely increased.
@@
Figure 1: (a) Schematic of sample synthesis procedure of sulfur doping in SCCLA. (b) XRD pattern of the S1 sample after 24 hours soaking in sulfide solution. (c) The morphological characteristics with the hexagonal grains highlighted. (d) Schematic of the directional stacking mechanism under high pressure.
@@
Pure copper-substituted lead apatite was first synthesized through a hydrothermal-calcination method following the procedure outlined in the literature Lee et al. (2023a); Wang et al. (2024), with the resulting gray powder being insulating. In order to carry out sulfur doping, as sketched in Fig. 1(a), the powder is ball-milled to nano-scaled particles and then thoroughly dispersed in a saturated solution of sodium sulfide. After stirring for 12 hours, the sample undergoes a 24-hour reaction in a high-pressure reaction vessel. The internal pressure of the reaction vessel is maintained at no less than 4 MPa, and the reaction temperature is set to 170 C. After the reaction, the powder is washed, filtered, dried, and then subject to a pressure of 20 MPa to compress it into a solid tablet. The resulting samples labeled as S1 and S2 are parallel with the latter was purified more carefully.
@@
The XRD spectra of S1 sample are refined using Reflex as shown in Fig. 1(b). Pawley Fitting is utilized to accurately determine lattice constants, and experimental and sample-related parameters. During the sulfur doping process, atoms within the apatite structure may have been etched or consumed, so the lattice is largely shrunk leading to a strong lattice distortion and good conductivity. The experimental spectra are consistent with that of the phosphate of P63m structure. By comparing to the standard of lead phosphate, the intensity differences observed at some peak positions could be attributed to the orientation of nano-scaled grains during growth in the sulfide solution. SEM was utilized to observe the morphological characteristics of the sample, as displayed in Fig. 1(c). The powder mainly consists of hexagonal flakes or hexagonal crystals. A small amount of hexagonal prism rod-like particle shows up, pointing out a significant difference compared to the sample before sulfur doping. This observation together with the analysis results of XRD spectra reveal the microstructure of the powder emerging as hexagonal grains, so the directional stacking is evident to function during high-pressure agglomeration, as sketched in Fig. 1(d). The head-to-tail quasi-1D stacking structure is then changed to be side-by-side, so the conducting channel might be broadened.
@@
Figure 2: (a) MT curves of ZFC and FC measurements of S1 sample at 25 and 200 Oe, respectively. The signal of an empty capsule is shown for comparisons. (b)-(e) MH hysteresis curves at eight temperatures. At 10 and 50 K, it is paramagnetic at strong field, and at higher temperature it becomes diamagnetic. Below 250 K, a remarkable magnetic hysteresis effect is observed between $-300$ and 300 Oe. At 300 and 350 K, although we have decreased the measurement interval, there is not visible hysteresis indicating the critical temperature is below 300 K. (f) MH hysteresis curve at 250 K is again finely measured with a quartz holder. The slight difference of lineshape may be due to the orientational difference of tablet samples on different holders.
@@
We then employed MPMS-3 SQUID to measure the dc magnetic moments of S1 sample, as shown in Fig. 2. It is clear from the magnetic moment–temperature (MT) curves, considering the background of capsule, the magnetization under magnetic field of 25 and 200 Oe is negative up to near room temperature, indicating its diamagnetism at weak field. The bifurcation between zero-field-cooling (ZFC) and field-cooling (FC) measurements is also found to be above 250 K. The moment–magnetic field (MH) curves at various temperatures from 10 to 350 K are also displayed. One can see that, from $-300$ to 300 Oe, the MH curves exhibit notable hysteresis effect up to 250 K. In particular, the hysteresis at 150 K is pretty obvious, exceeding the highest critical temperature of known superconductors at ambient pressure. This phenomenon has been reported in our previous paper Wang et al. (2024), but the quality of the present data is largely increased, further eliminating the possibility of measurement faults. At 300 and 350 K, the hysteresis can not be detected. As stated before, these results strongly suggest there exists a Meissner effect in SCCLA.
@@
Figure 3: (a) The RT curves and their derivative at current of 8 mA during heating and cooling the S1 sample. Based on the RT relationship, the temperature region can be divided into three: Fermi liquid, strange metal and bad metal. (b) RT curves of S2 sample at various magnetic field.
@@
Measurement results of transport with indium electrode are displayed in Fig. 3, which are mainly conducted on PPMS and have been reconfirmed on other facilities. Although the sample before soaking in sulfur is insulator, after soaking the resistivity at room temperature is largely reduced to around $2\times 10^{-5}~{}\Omega\cdot$m, close to that of natural graphite. It is thus evident that the sulfur-copper codoping plays essential role in the improvement of transport properties of the insulating apatite ionic crystal. Such considerable conductivity is far beyond expectation, that has to be carefully comprehended.
@@
As estimated by the critical magnetic field, if there is a superconducting phase, the critical current should be as small as $\mu$A, which can not be accurately detected in our present facility. The RT curves are therefore measured with constant current of 8 mA to ensure the precision, as illustrated in Fig. 3. It is found that below 20 K the derivative $dR/dT$ exhibits a nearly linear relation suggesting a Fermi liquid behavior as in normal metals. Increasing the temperature, the derivative becomes nearly constant revealing a linear-in-temperature resistivity, which is an essential signature of strange metal. At around 230 K in the cooling curve and 304K in the heating curve, there are remarkable turning points to figure out the transition from strange metal to bad metal phase. The transition is more evident in S2 sample than that in S1, implying it is a phase transition of second order. It has been stated that, if the resistivity in bad metal can be further reduced by one or two orders Yang et al. (2022), these transition points can then be recognized as the superconducting critical temperature. The thermal hysteresis of the two RT curves, probably stemming from the poor thermal conductivity of the samples, is also consistent with that of the magnetic measurements.
Figure 4: (a) MT curves of S3 sample with quartz holder at 25 and 200 Oe for ZFC and FC measurements. Inset displays the MT curves at 25 Oe with capsule. The bifurcation occurs around 250 K. (b) MH curves at three typical temperatures. The diamagnetism is too strong at 10 K, so it is also displayed with larger scale in the inset. (c) RT curves at 100 $\mu$A with the electrodes being silver and indium, respectively. Inset shows the temperature range within 0 to 150 K. At low temperature, the resistance of silver electrode sample is too small to be detectable due to the limit of instrument.
[…]
The RT curves of the same sample are individually measured with different electrodes, silver and indium. It is quite strange that, with silver the room-temperature resistance is much larger than that with indium, but at low temperature, the resistance with silver becomes much smaller, even undetectable. The transition point with silver electrode is also around 250 K, consistent with that of diamagnetism, excluding the possibility of virtual connection. It is noticed that, after measurement, the color of silver electrode becomes black implying it has been sulfurized. We therefore think that, the silver extracts the sulfur in the 1D ionic channel of apatite, leading to an electron-rich state and thus strengthening the superconductivity. The indium electrode can however not closely contact to the apatite structure, so it solely performs a strange-metal behavior.
@@
Compared to our previous synthetic procedure Wang et al. (2024), the doping of sulfur not only facilitates the performance of copper, but also changes the topology of 1D ionic channel of apatite. From head-to-tail to side-by-side, the quasi-1D copper-sulfur lattice in apatite may have got much stronger inter-chain interaction to activate tunneling in between. The alkali metals may also play roles in connecting the crystal grains. This explains the huge improvement of electric conductivity and the emergence of Meissner effect. We have also conducted microwave absorption measurement and did not find visible radical signals, implying the absence of quasiparticles in normal metal. People might suspect the magnetic hysteresis is enabled by some pseudogap states, but since the RT curves at least exhibit a metallic feature and extremely low resistance in lead-free samples, we would be more preferable to think it is because the superconducting critical current is below our measurement limit. During measurements, we have noticed that the SCCLA sample possesses abnormally large electric capacity and poor thermal conductivity Bi et al. (2022); Deppe et al. (2004), which are inconsistent with the high electric conductivity. In particular, the capacity often induces strange sudden change between low- and high-resistance states enabling large fluctuations of RT curves, even if we have carefully excluded other interferences. It is more likely the sample is continuously charging and discharging during electric measurements, hindering the detectability of zero resistance in SCCLA. Considering the performance of lead-free sample, the dielectrics of lead apatite backbone matters.
@@
In summary, we modify the synthetic procedure of SCCLA to codope both sulfur and copper into lead apatite, and the structural characterization reveals a directional stacking mechanism. The magnetic and electric properties of SCCLA have been comprehensively investigated. The hysteresis MH loops can be observed up to 250 K, and the ZFC–FC bifurcation occurs as well. The RT curve manifests that SCCLA possesses a strange-metal phase at large current and a second-order phase transition occurs at around 230 K during cooling. Further synthesis of lead-free sample is then performed which provides us even stronger diamagnetism and smaller resistance at low temperature. We therefore believe that we have made a substantial step towards room-temperature superconductivity.